Please fully explain

1) Consider two charged particles (q,--4μC,m1-8 grams; q,-+15pC, m2-3 grams). The first particle is 18 cm to the North of the second one. a) What is the magnitude and direction of the electric FORCE on the 1st particle due to the presence of the 2nd particle? [when answering this direction question, answer with North, East, South, West, Up or Down.] 16.7 Nt South b) Is the magnitude of the force on the 2nd particle due to the 1st particle bigger, the same as, or smaller than the force on the 1st particle due to the 2nd particle? same c) Is the direction of the force on the 1st particle due to the 2nd particle the same or opposite that of the force on the 2nd particle due to the 1st particle? opposite d) Is the magnitude of the acceleration of the 2nd particle due to the presence of the 1st particle bigger, the same as, or smaller than the acceleration of the 1st particle due to the 2nd particle? (Assume there are no other forces acting on the two particles) bigger 1st particle <-W E> 18 cm +. 2d particle
Answer #1

Like charges repel unlike charges repel . electro static force is opposite and equal between the charges .charge a exerts equal and opposite force on charge B as charge b exerts on charge a .

F= kq1q2/r^2= 8.99 * 10^9 * 15* 4* (10^-12) / (18*10^-2)^2 = 1.66 * 10 =16.6 N

2nd charge positive attracts 1st particle with a force equal to F that is south

magnitude of force will be equal but in opposite in direction for charges each one attract in its direction.

acceleration = F/m so it is inversely proportional to mass smaller mass more acceleration 2nd particle mass is smaller compared to 1st so 2nd particle will have bigger acceleration .